A ball of mass 0.3 kg is released from a tower of height 100 m, simultaneously a bullet of mass 0.2 kg fired vertically upward from point just below it with speed 100 m/s find height attained above the tower by combined mass system after collision 

Grade:11

1 Answers

Arun
5 years ago
Dear Sumit
 
 
velocity of combined mass = 0.5*v = 0.3*0 + 0.2*100 = 40 m/s
height attained above the tower = u2/2g = 1600/20 = 80m
 
Hope it helps
 
Regards
Arun